I need to practice these two question, please I need a good explanation of these exercises.

Many Thanks

The following data exists for Shop Rite Corporation:

Beginning of Year   End of Year

Total Assets $600     $1,200

Total Liabilities $300 $500

During the year, the company issued $500 stock and paid $300 in dividends to its existing stockholders. How much was net income or loss for the year?

On April 1, a consulting company has total assets of $60,000. It then has 4 transactions as follows:

          i.              Earned $20,000 revenue for services completed. Client promises to pay in 30 days.

        ii.              The same client makes a payment of $12,000 bringing their balance down to $8,000.

      iii.              Make a cash payment on a note payable of $6,000.

       iv.              Buy a $25,000 building for $12,000 in cash and $13,000 in a note payable.

How much are total assets after these transactions?
Answer #1

1) Beginning Stockholder's equity = 600-300 = 300

Ending Stockholder's equity = 1200-500 = 700

Net income = Beginning equity+Stock issued-Dividend paid-Ending equity = 300+500-300-700 = 200

2) Calculate total assets

Beginning balance 60000
i) 20000
ii) No effect
iii) -6000
iv) 13000
Total Assets 87000